OpenStudy (anonymous):

What is an equation of a line passing through (3,3) and (4,7) show working

OpenStudy (anonymous):

Again, first find the slope using (y2-y1)/(x2-x1)

OpenStudy (anonymous):

@Mhansen let x1= 3, y1 = 3, x2 = 4, y2 = 7

OpenStudy (anonymous):

(7-3)/(4-3)

OpenStudy (anonymous):

Yes I did that

OpenStudy (anonymous):

what is the slope? Put that into point-intercept form, which is: y-y1=m(x-x1), where m is the slope, and x1/y1 is the x and y coordinate of one of the pairs of points given.

OpenStudy (anonymous):

Lets use the point (3,3). y-(3)=4(x-3) solve for y

OpenStudy (anonymous):

*distribute the 4 first, though

OpenStudy (anonymous):

Oh thx

OpenStudy (anonymous):

post what you got

OpenStudy (anonymous):

It should be in the form of y=mx+b

OpenStudy (anonymous):

Having trouble?

OpenStudy (anonymous):

Yes